# Reason2Decide: Rationale-Driven Multi-Task Learning

### v1

@@ Abstract
Despite the wide adoption of Large Language Models (LLM)s, clinical decision support systems face a critical challenge: achieving high predictive accuracy while generating explanations aligned with the predictions. Current approaches suffer from exposure bias leading to misaligned explanations. We propose Reason2Decide, a two-stage training framework that addresses key challenges in self-rationalization, including exposure bias and task separation. In Stage-1, our model is trained on rationale generation, while in Stage-2, we jointly train on label prediction and rationale generation, applying scheduled sampling to gradually transition from conditioning on gold labels to model predictions. We evaluate Reason2Decide on three medical datasets, including a proprietary triage dataset and public biomedical QA datasets. Across model sizes, Reason2Decide outperforms other fine-tuning baselines and some zero-shot LLMs in prediction (F1) and rationale fidelity (BERTScore, BLEU, LLM-as-a-Judge). In triage, Reason2Decide is rationale source-robust across LLM-generated, nurse-authored, and nurse-post-processed rationales. In our experiments, while using only LLM-generated rationales in Stage-1, Reason2Decide outperforms other fine-tuning variants. This indicates that LLM-generated rationales are suitable for pretraining models, reducing reliance on human annotations. Remarkably, Reason2Decide achieves these gains with models $40\times$ smaller than contemporary foundation models, making clinical reasoning more accessible for resource-constrained deployments while still providing explainable decision support.
@@
H M Quamran Hasan, Housam Khalifa Bashier, Jiayi Dai, Mi-Young Kim, Randy Goebel
University of Alberta
Edmonton, Alberta, Canada
 Camrose, Alberta, Canada
@@ 1.  Introduction
The integration of reasoning capabilities with prediction tasks has been a critical research problem in natural language processing (NLP). Existing state-of-the-art language models struggle to balance high predictive accuracy while also generating human-interpretable explanations niu-etal-2025-knowledge. Although LLMs have demonstrated strong performance on various question-answering benchmarks hendrycks2021measuringmassivemultitasklanguage; srivastava2023imitationgamequantifyingextrapolating, their ability to provide rationales that align with their predictions remains limited. This limitation is substantially amplified in healthcare where explanations are essential for trust and adoption.
Current approaches for rationale generation face two fundamental challenges. First, the exposure bias problem schmidt-2019-generalization, which arises when models are trained to generate rationales conditioned only on ground-truth labels, but need to explain their own potentially incorrect predictions during inference. Second, the task separation problem, which occurs when prediction and rationale generation are treated independently rather than as mutually supporting tasks narang2020wt5trainingtexttotextmodels. While multi-task learning has shown promising performance, existing methods have failed to address the train-test discrepancy: models exclusively learn to explain gold-standard labels during training, leaving them unprepared for rationalizing their own prediction during inference.
Recent work has acknowledged these challenges and explored various approaches to improve model interpretability. For example, “Chain-of-Thought” (CoT) prompting NEURIPS2022_9d560961 demonstrates that step-by-step reasoning can improve complex reasoning; similarly, self-consistency techniques wang2023selfconsistencyimproveschainthought show the value of aggregating multiple reasoning paths. However, these methods primarily focus on inference time strategies rather than addressing the fundamental training dynamics that help establish more robust reasoning capabilities. Knowledge distillation approaches hsieh-etal-2023-distilling and rationale-augmented training lampinen2022tellwhyexplanationssupport have also shown benefits, but do not train models to generate explanations conditioned on their own predictions during training.
Here we propose Reason2Decide (Figure 1), a framework which trains a single model to jointly predict and generate rationales. Our approach consists of a two-stage training regime, based on insights from curriculum learning 10.1145/1553374.1553380 and scheduled sampling 10.5555/2969239.2969370. The idea is to first learn to model explanation fundamentals and then jointly produce predictions and rationales, where we gradually shift from gold-label conditioning to self-conditioning. Through this process, the model learns to explain its own predictions. This addresses the exposure bias problem while promoting alignment between predictions and rationales.
@@ 2.1.  Rationale Generation and Explainable AI
A primary goal of Explainable AI (XAI) is making model decisions transparent and interpretable. The development of early XAI methods focused on post-hoc explanation generation by analyzing model internals or feature importance ribeiro2016whyitrustyou; lundberg2017unifiedapproachinterpretingmodel. But recently, the field has shifted towards self-explaining models that can learn to generate rationales simultaneously with their predictions camburu2018esnlinaturallanguageinference. Our work aligns with the latter. For instance, narang2020wt5trainingtexttotextmodels fine-tuned T5 raffel2023exploringlimitstransferlearning models to generate explanations by treating them as a separate task. Similarly, lampinen2022tellwhyexplanationssupport demonstrate that incorporating natural language explanations during training can improve reasoning and generalization. However, these approaches often treat the explanation as a secondary output, which creates a separation between the prediction and its explanation. In contrast, our Reason2Decide framework is designed to align the predictive task with the explanatory one, to help ensure that the explanation is an integral part of the reasoning process.
@@
Multi-Task Learning aims to improve generalization by leveraging shared representations across related tasks ruder2017overviewmultitasklearningdeep; 10.1023/A:1007379606734. In NLP, models like T5 are considered multi-task learners by combining diverse problems into a unified text-to-text format. We similarly use this foundation but specifically focus on the alignment between prediction and rationales. Closely related work includes knowledge distillation approaches that use rationales. hsieh-etal-2023-distilling distills the capabilities of a larger teacher model that generates step-by-step rationales into a smaller student model. While this approach can be effective in improving predictive accuracy, it does not explicitly enforce prediction–explanation coordination. The student model benefits from teacher rationales but is not trained to ensure that its own predictions and explanations are self-consistent. Our method addresses this by explicitly training the coordination between prediction and rationale generation.
@@ 2.3.  Mitigating Exposure Bias and Scheduled Sampling
Exposure bias in self-rationalization occurs when models are trained to generate explanations conditioned on gold labels but at inference time must justify their own predictions. Scheduled sampling 10.5555/2969239.2969370 addresses the issue by gradual replacement of teacher-forced tokens with model-generated ones. Here, we adapt this principle to the task level, thus treating the model’s predicted label as the conditioning context for rationale generation. Our mechanism systematically transitions from using gold labels to the model’s own predictions, directly mitigating this bias and preparing models for real-world deployment.
@@ 3.1.  Stage-1: Rationale Foundation Training
Previous work has shown that domain-specific pretraining significantly improves model performance in specialized tasks lewis-etal-2020-pretrained; 10.1093/bioinformatics/btz682. With this motivating background, our first stage teaches explanation fundamentals by training the model to generate rationales. Let $\theta$ denote the parameters of the T5 model. Given a gold rationale $r^{*}$, we optimize:
@@ 3.2.  Stage-2: Joint optimization of Prediction and Explanation with Task-Level Scheduled Sampling
To mitigate exposure bias in self-rationalization, we introduce a scheduled sampling mechanism. Unlike prior token-level scheduled sampling 10.5555/2969239.2969370, which replaces target tokens within sequences, we apply it at the task level by switching the conditioning context from gold labels to model-predicted labels. For instance, instead of gradually replacing reference tokens in a generated sequence, our method gradually replaces reference labels used for conditioning the rationale generation.
@@
The ceiling $0.9$ was a design choice to prevent the model from fully relying on self-generated labels during training to help avoid error amplification from incorrect predictions. The hyperparameters ($0.05$ and $0.6$) were determined through hyperparameter optimization using a few sets of predefined tuples, and are motivated by two key considerations:
@@
During $0<t<w$, we linearly increase $\alpha_{t}$ from 0 to $0.7$. This is to prioritize explanation loss $\mathcal{L}_{\text{expl}}$ over prediction loss $\mathcal{L}_{\text{pred}}$, to provide a smoother transition from single-task to multi-task optimization. After this phase $\alpha_{t}$ is held constant at $0.7$ for the rest of the training. We heuristically selected $0.7$ to balance both tasks, with a slight bias towards predictions, as Stage-1 already teaches rationale generation.
@@
During $w\leq t<w+m$, we gradually increase the fraction of predicted labels used for conditioning from 0 to $0.9$. This transition period is to systematically address exposure bias by training the model to learn coherent rationale generation conditioned on its own (potentially imperfect) predictions, bridging the gap between training and inference conditions.
@@ 4.  Experiments
In this section we introduce the datasets, provide implementation details, followed by the experimental results.
@@ 4.1.1.  Clinical Triage Dataset
The Clinical Triage Dataset is derived from a Canadian provincial health service. The dataset contains triage notes authored by nurses based on telephone triage dialogue with calling patients, along with the recommended care disposition and a rationale for that disposition choice. Specific data source details are omitted to preserve anonymity.
Task: Given a nurse triage note, predict a disposition (care pathway) and generate a rationale for the decision. The label space consists of 12 classes ranging from low (Home Care) to high urgency (Go to Emergency Department (ED) now). [^1] A complete list of the 12 disposition classes is provided in the Appendix, Subsection 9.2
@@
Table 1 shows sample rationales of each type. The post-processed version is there because the original version consists of atomic facts and are not always grammatically correct. We use  [^2] All prompts used are provided in the Appendix, Subsection 9.1. a Qwen-3 8B qwen3technicalreport model to convert them into grammatically correct full sentences without adding new information. For LLM-generated notes, we provide the model the triage note and disposition and ask it to explain why the disposition was chosen. For experiments on the whole dataset, we used the LLM-generated rationales.
@@
Data split: The train, validation and test sets consist of 170,695, 21,312 and 9,704 samples respectively.
@@ 4.1.2.  PubMedQA
Task: Biomedical question answering with Yes/No/Maybe labels jin-etal-2019-pubmedqa. We concatenate the question and context as the model input as triage note replacement.
@@ 4.1.3.  BioASQ (Task B, Yes/No)
Task: Biomedical question answering dataset from the BioASQ 13 challenge nentidis2025overviewbioasq2025thirteenth. For our experiments, we focus on yesno questions (binary classification).
Rationales: We extract and concatenate the snippets to serve as rationales for each question. Because the concatenated snippets can be long, we first summarize them with a Qwen-3-8B model and use the summary as the gold rationale.
Data split: The official training dataset consists of 1,459 yesno questions. We randomly select 250 questions for our validation set, and 1,209 questions for the train set. For the test set, we use the concatenation of the four official test batches, yielding 82 questions.
@@ 4.2.  Implementation Details
We implement Reason2Decide using T5 Small (77M), Base (250M), Large (800M) architectures using A100 X 4 GPUs. [^3] Source code available at: https://github.com/quamranhasan/Reason2Decide Training uses AdamW optimizer loshchilov2019decoupledweightdecayregularization with learning rate $5\times 10^{-5}$, max input length = 1024, and effective batch size = 64. Model-specific configuration is provided in Table 2.
@@ 4.3.  Baselines
  • Standard fine-tuning (SFT): Single-task label prediction without rationales howard-ruder-2018-universal; model selection by validation Macro-F1.
  • Distilling Step-by-Step (DSS): Multi-task training following hsieh-etal-2023-distilling. We run two selection criteria:
    • DSS-Loss: Model selection by validation loss (as in hsieh-etal-2023-distilling).
@@
  • Zero-shot LLMs: Zero-shot baselines without task-specific fine-tuning, using open-source models including Qwen-3-8B and Qwen-3-32B qwen3technicalreport, as well as Llama-3.1-Aloe-Beta-8B garcia2025aloe.
@@
For all fine-tuned baselines, we use the same optimizer, learning rate, effective batch size as our method; per-device batch sizes are set following Table 2. Early stopping is used with patience of 5 validation evaluations. Model selection is by validation Macro-F1 unless otherwise stated (DSS-Loss). We report means over three seeds/runs. We use greedy decoding to ensure run-to-run determinism.
@@
  • BERTScore zhang2020bertscoreevaluatingtextgeneration (F1 variant): Semantic similarity between generated and gold rationales.
  • BLEU papineni-etal-2002-bleu: N-gram overlap for surface-level quality.
  • LLM-as-a-Judge: Following recent work demonstrating strong alignment between LLM and human evaluation zheng2023judgingllmasajudgemtbenchchatbot; niu-etal-2025-knowledge, we employ Qwen-3-8B for expert-style assessment on random 2k samples for the Triage Dataset and PubMedQA, and the whole test set for BioASQ (82 samples). We evaluate three defined metrics:

### v2

@@ Abstract
Despite the wide adoption of Large Language Models (LLM)s, clinical decision support systems face a critical challenge: achieving high predictive accuracy while generating explanations aligned with those predictions. Current approaches suffer from exposure bias, leading to misaligned explanations. We propose Reason2Decide, a two-stage training framework that addresses key challenges in self-rationalization, including exposure bias and task separation. In Stage-1, our model is trained on rationale generation, while in Stage-2, we jointly train on label prediction and rationale generation, applying scheduled sampling to gradually transition from conditioning on gold labels to model predictions. We evaluate Reason2Decide on three medical datasets, including a proprietary triage dataset and public biomedical QA datasets. Across model sizes, Reason2Decide outperforms other fine-tuned baselines and some zero-shot LLMs in prediction (F1) and rationale fidelity (BERTScore, BLEU, LLM-as-a-Judge). In triage, Reason2Decide is rationale source-robust across LLM-generated, nurse-authored, and nurse-post-processed rationales. In our experiments, while using only LLM-generated rationales in Stage-1, Reason2Decide outperforms other fine-tuned variants. This indicates that LLM-generated rationales are suitable for pretraining models, reducing reliance on human annotations. Remarkably, Reason2Decide achieves these gains with models $40\times$ smaller than contemporary foundation models, making clinical reasoning more accessible for resource-constrained deployments while still providing explainable decision support.
@@
H M Quamran Hasan1, Housam Khalifa Bashier1, Jiayi Dai1,
Mi-Young Kim2, Randy Goebel1
1University of Alberta, Edmonton, Alberta, Canada
2University of Alberta, Camrose, Alberta, Canada
@@ 1.  Introduction
The integration of reasoning capabilities with prediction tasks has been a critical research problem in natural language processing (NLP). Existing state-of-the-art large language models (LLMs) struggle to balance high predictive accuracy while also generating human-interpretable explanations Niu et al. (2025). Although LLMs have demonstrated strong performance on various question-answering benchmarks Hendrycks et al. (2021); Srivastava and others (2023), their ability to provide rationales that align with their predictions remains limited. This limitation is substantially amplified in healthcare where explanations are essential for trust and adoption.
Current approaches for rationale generation face two fundamental challenges. First, the exposure bias problem Schmidt (2019), which arises when models are trained to generate rationales conditioned only on ground-truth labels, but need to explain their own potentially incorrect predictions during inference. Second, the task separation problem, which occurs when prediction and rationale generation are treated independently rather than as mutually supporting tasks Narang et al. (2020). While multi-task learning has shown promising performance, existing methods have failed to address the train-test discrepancy: models exclusively learn to explain gold-standard labels during training, leaving them unprepared for rationalizing their own predictions during inference.
Recent work has acknowledged these challenges and explored various approaches to improve model interpretability. For example, “Chain-of-Thought” (CoT) prompting Wei et al. (2022) demonstrates that step-by-step reasoning can improve complex reasoning; similarly, self-consistency techniques Wang et al. (2023) show the value of aggregating multiple reasoning paths. However, these methods primarily focus on inference time strategies rather than addressing the fundamental training dynamics that help establish more robust reasoning capabilities. Knowledge distillation approaches Hsieh et al. (2023) and rationale-augmented training Lampinen et al. (2022) have also shown benefits, but do not train models to generate explanations conditioned on their own predictions during training.
Here we propose Reason2Decide (Figure 1), a framework which trains a single model to jointly predict and generate rationales. Our approach consists of a two-stage training regime, based on insights from curriculum learning Bengio et al. (2009) and scheduled sampling Bengio et al. (2015). The idea is to first learn to model explanation fundamentals and then jointly produce predictions and rationales, where we gradually shift from gold-label conditioning to self-conditioning. Through this process, the model learns to explain its own predictions. This addresses the exposure bias problem while promoting alignment between predictions and rationales.
@@ 2.1.  Rationale Generation and Explainable AI
A primary goal of Explainable AI (XAI) is making model predictions transparent and interpretable. The development of early XAI methods focused on post-hoc explanation generation by analyzing learned model internal attributes or feature importance Ribeiro et al. (2016); Lundberg and Lee (2017). Recently, the field has shifted towards self-explaining models that can learn to generate rationales simultaneously with their predictions Camburu et al. (2018). Our work aligns with the latter. For instance, Narang et al. (2020) fine-tuned T5 Raffel et al. (2023) models to generate explanations by treating them as a separate task. Similarly, Lampinen et al. (2022) demonstrate that incorporating natural language explanations during training can improve reasoning and generalization. However, these approaches often treat the explanation as a secondary output, which creates a separation between the prediction and its explanation. In contrast, our Reason2Decide framework is designed to align the predictive task with the explanatory one, to help ensure that the explanation is an integral part of the reasoning process.
@@
Multi-Task Learning aims to improve generalization by leveraging shared representations across related tasks Ruder (2017); Caruana (1997). In NLP, models like T5 are considered multi-task learners because they combine diverse problems into a unified text-to-text format. We similarly use this foundation but specifically focus on the alignment between prediction and rationales. Closely related work includes knowledge distillation approaches that use rationales. For example, Hsieh et al. (2023) distills the capabilities of a larger teacher model that generates step-by-step rationales into a smaller student model. While this approach can be effective in improving predictive accuracy, it does not explicitly enforce prediction–explanation coordination. The student model benefits from teacher rationales but is not trained to ensure that its own predictions and explanations are self-consistent. Our method addresses this weakness by explicitly training the coordination between prediction and rationale generation.
@@ 2.3.  Mitigating Exposure Bias and Scheduled Sampling
Exposure bias in self-rationalization occurs when models are trained to generate explanations conditioned on gold labels but at inference time must justify their own predictions. Scheduled sampling Bengio et al. (2015) addresses the issue by gradual replacement of teacher-forced tokens with model-generated ones. Here, we adapt this principle to the task level, thus treating the model’s predicted label as the conditioning context for rationale generation. Our mechanism systematically transitions from using gold labels to the model’s own predictions, directly mitigating this bias and preparing models for real-world deployment.
@@ 3.  Methodology
We employ a single encoder-decoder architecture $f_{\theta}$ (T5 variants) without modifications to its core components including the number of attention heads, hidden layer dimensions, relative positional encodings, and activation functions which are identical to the original specification. T5’s text-to-text formulation is ideal for our rationale-driven paradigm, as it naturally supports autoregressive decoding for both discrete label prediction (e.g., “Yes", “No", or “Go to ED Now.") and free-form rationale generation from task-specific prompts. This allows seamless multi-tasking with explanations under a single objective, unlike encoder-only models like BERT Devlin et al. (2019), which require separate classification heads and lack native support for conditional explanation generation.
@@ 3.1.  Stage-1: Rationale Foundation Training
Previous work has shown that domain-specific pretraining significantly improves model performance in specialized tasks Lewis et al. (2020); Lee et al. (2019). With this motivating background, our first stage teaches explanation fundamentals by training the model to generate rationales. Let $\theta$ denote the parameters of the T5 model. Given a gold rationale $r^{*}$, we optimize:
@@ 3.2.  Stage-2: Joint Optimization of Prediction and Explanation with Task-Level Scheduled Sampling
To mitigate exposure bias in self-rationalization, we introduce a scheduled sampling mechanism. Unlike prior token-level scheduled sampling Bengio et al. (2015), which replaces target tokens within sequences, we apply it at the task level by switching the conditioning context from gold labels to model-predicted labels. For instance, instead of gradually replacing reference tokens in a generated sequence, our method gradually replaces reference labels used for conditioning the rationale generation.
@@
The ceiling of $0.9$ was a design choice to prevent the model from fully relying on self-generated labels during training, in order to help avoid error amplification from incorrect predictions. The hyperparameters ($0.05$ and $0.6$) were determined through hyperparameter optimization using a few sets of predefined tuples, and are motivated by two key considerations:
@@
During $0<=t<w$, we linearly increase $\alpha_{t}$ from 0 to $0.7$. This is to prioritize explanation loss $\mathcal{L}_{\text{expl}}$ over prediction loss $\mathcal{L}_{\text{pred}}$, to provide a smoother transition from single-task to multi-task optimization. After this phase $\alpha_{t}$ is held constant at $0.7$ for the rest of the training. We heuristically selected $0.7$ to balance both tasks, with a slight bias towards predictions, as Stage-1 already teaches rationale generation.
@@
Exposure bias arises in autoregressive rationale generation from the train-test mismatch: training conditions explanations on gold labels, but inference must use model predictions, risking error propagation. During $w\leq t<w+m$, we gradually increase the fraction of predicted labels used for conditioning from 0 to $0.9$. This transition period is to systematically address exposure bias by training the model to learn coherent rationale generation conditioned on its own (potentially imperfect) predictions, bridging the gap between training and inference conditions.
@@ 4.  Experiments
Here we introduce the datasets and provide implementation details, followed by the experimental results.
@@ 4.1.1.  Clinical Triage Dataset
The Clinical Triage Dataset is derived from Alberta Health Link 811, a provincial telephone-based health advice service operated by Alberta Health Services (AHS), Canada. The dataset contains triage notes authored by nurses during telephone consultations with patients, along with the recommended care disposition and a rationale explaining the decision.
Task: Given a nurse triage note, predict a disposition (care pathway) and generate a rationale for that recommendation. The label space consists of 12 classes ranging from low (Home Care) to high urgency (e.g., "Go to Emergency Department (ED) now"). [^1] A complete list of the 12 disposition classes is provided in the Appendix, Subsection 10.2
@@
Table 1 shows sample rationales of each type. The post-processed version is there because the original version consists of atomic facts and is not always grammatically correct. We use [^2] All prompts used are provided in the Appendix, Subsection  10.1. a Qwen-3 8B Team (2025) model to convert them into grammatically correct full sentences without adding new information. For LLM-generated notes, we provide the model the triage note and disposition and ask it to explain why the disposition was chosen. For experiments on the whole dataset, we used the LLM-generated rationales.
@@
Data split: The train, validation and test sets consist of roughly 171k, 21k and 9.7k samples, respectively.
@@ 4.1.2.  PubMedQA
Task: Biomedical question answering with Yes/No/Maybe labels Jin et al. (2019). We concatenate the question and context as the model input as triage note replacement.
@@ 4.1.3.  BioASQ (Task B, Yes/No)
Task:
Biomedical question answering dataset from the BioASQ 13 challenge Nentidis et al. (2025). For our experiments, we focus on yes/no questions (binary classification).
Rationales:
We extract and concatenate the snippets to serve as rationales for each question. Because the concatenated snippets can be long, we first summarize them with a Qwen-3-8B model and use the summary as the gold rationale.
Data split:
The official training dataset consists of 1,459 yes/no questions. We randomly select 250 questions for our validation set, and 1,209 questions for the train set. For the test set, we use the concatenation of the four official test batches, yielding 82 questions.
@@ 4.2.  Implementation Details
We implement [^3] Source code available at: https://github.com/quamranhasan/Reason2Decide Reason2Decide using T5 Small (77M), Base (250M), Large (800M) architectures using A100 X 4 GPUs. Training uses AdamW optimizer Loshchilov and Hutter (2019) with learning rate $5\times 10^{-5}$, max input length = 1024, and effective batch size = 64. Model-specific configuration is provided in Table 2.
@@ 4.3.  Baselines
  • Standard fine-tuning (SFT): Single-task label prediction without rationales Howard and Ruder (2018); model selection by validation Macro-F1.
  • Distilling Step-by-Step (DSS): Multi-task training following Hsieh et al. (2023). We run two selection criteria:
    • DSS-Loss: Model selection by validation loss (as in Hsieh et al. (2023)).
@@
  • Zero-shot LLMs: Zero-shot baselines without task-specific fine-tuning, using open-source models including Qwen-3-8B and Qwen-3-32B Team (2025), as well as Llama-3.1-Aloe-Beta-8B Garcia-Gasulla et al. (2025).
@@
For all fine-tuned baselines, we use the same optimizer, learning rate, and effective batch size as in our method; per-device batch sizes are set following Table 2. Early stopping is used with patience of 5 validation evaluations. Model selection is by validation Macro-F1 unless otherwise stated (DSS-Loss). We report means over three seeds/runs. We use greedy decoding to ensure run-to-run determinism.
@@
  • BERTScore Zhang et al. (2020) (F1 variant): Semantic similarity between generated and gold rationales.
  • BLEU Papineni et al. (2002): N-gram overlap for surface-level quality.
  • LLM-as-a-Judge: Following recent work demonstrating strong alignment between LLM and human evaluation Zheng et al. (2023); Niu et al. (2025), we employ Qwen-3-8B for expert-style assessment on random 2k samples for the Triage Dataset and PubMedQA, and the whole test set for BioASQ (82 samples). We evaluate three defined metrics: